OpenStudy (anonymous):

sequence and series: find the fraction form of 3.02. note: 2 is recuring

OpenStudy (anonymous):

151/50

OpenStudy (anonymous):

how do you get that from

jimthompson5910 (jim_thompson5910):

Let x = 3.0222222... So 10x = 30.222 and 100x = 302.22222... which means that 100x - 10x = 302.2222... - 30.2222... 90x = 272 x = 272/90 x = 136/45 So 136/45 = 3.022222...

OpenStudy (anonymous):

I used aTI-84 calculator

OpenStudy (anonymous):

I need to use this formula S=a1/1-r